A golf hole is exactly in the middle of a circular green. If the distance from the center of the hole to the edge of the green is 25 feet, what is the area of the green?
Answer by checkley75(3666)

A=PIR^2
A=22/7*25^2
A=22/7*625
A=13750/7
A=1964.2857 SQUARE FEET FOR THE AREA OF THE GREEN